I need help to label the HNMR of this compound. Can someone help? Can you tell me what each peak mean? For example, 2H triplet

Diels Alder Adduct Lactone HNMR 07/23/2020 CO2H PER 0.07 0.98 09 130 1.05 1.07 3.06 1.26
Answer #1

Solution:- OH 1 HNMR a = 1.11, (d, 3H) b = 2.50, (m. 1H) g H fh f c=2.98, (dd, 1H) d! d=3.18, (m, 1H) h b e=3.38, (dd, 1H) OH